OpenStudy (anonymous):

What is the standard form equation of the line passing through (–2, 8) with a slope of 2?

OpenStudy (anonymous):

can someone explain to me how to do this step by step

OpenStudy (anonymous):

what is the formula i will be using for this

OpenStudy (anonymous):

y=mx+b Where M is the slope and b is they y intercept

OpenStudy (anonymous):

8=2(-2)+b

OpenStudy (anonymous):

is this right so far?

OpenStudy (anonymous):

8=-4+b

OpenStudy (anonymous):

Looks right if you don't have the y intercept.

OpenStudy (anonymous):

how do i find the y intercept

OpenStudy (anonymous):

Hello there. The standard line form is something like this: Ax + By = C or Y = mx + C where m is the slope. So, if you put the values in, you will have something like 8 = -2x + C For more information, you can look this video up: http://www.youtube.com/watch?v=-8yM86_1fC4 Sorry, I'd help but they will ban me for being "too direct".

OpenStudy (anonymous):

Take x as 0 in the equation to find the y-intercept.

OpenStudy (anonymous):

i dont understand what you mean @Fall12-13

OpenStudy (anonymous):

Solve for b. and that's your y intercept. if the two points you solved have the slope 2. You have the right answer.

OpenStudy (anonymous):

b=4?

OpenStudy (anonymous):

@Nikee We cannot be too direct on this one.

OpenStudy (anonymous):

2x – y = 12 2x – y = –12 2x + y = 12 2x + y = –12

OpenStudy (anonymous):

these are my answers

OpenStudy (anonymous):

i think the answer is b? is that right @Sir_Rico_of_Eureka

OpenStudy (lstrasmann):

Is it multiple choice? If it is can you give us the choices?

OpenStudy (anonymous):

2x – y = 12 2x – y = –12 2x + y = 12 2x + y = –12

OpenStudy (anonymous):

i think its B but i dont know this is my last question on my pretest and i want to have all the correct answers

OpenStudy (anonymous):

hello???

OpenStudy (lstrasmann):

Im thinking about it. Give me a second.

OpenStudy (anonymous):

did u figure it out yet

OpenStudy (lstrasmann):

I think you have to plug in the y-intercept (i found out it is 8 + 2x) into\[8 = -2x + C \]like this\[8 = -2x + 8+2x\]then subtract 8\[0 = -2x + 2x\]then add -2x and 2x\[0=0\]wait, maybe that's wrong. but that's what i would do, if it isn't a choice i would just guess at that point.

OpenStudy (anonymous):

Sorry I had to get to class. If you plugged it all in and got 4 for b then you'll get y=2x+4 which is 2x-y =-4 class is starting but redo your math

OpenStudy (anonymous):

2*-2 = -4